Write a polynomial function of least degree with integral coefficients whose zeros include 4 and 2i.
Answer by fractalier(2101)

Since 2i is a root, so must be -2i. Thus our polynomial starts as
(x - 2i)(x + 2i)(x - 4) = 0
(x^2 + 4)(x - 4) = 0
x^3 - 4x^2 + 4x - 16 = 0